Development of a Computer-Assisted Tool for Morphological Studies

Haruyuki Tatsumi, Masahiro Nakamura

Department of Anatomy, Sapporo Medical University School of Medicine, Sapporo, Japan E-mail:tatsumi@sapmed.ac.jp, nakamura@sapmed.ac.jp

1 Introduction

In mophological studies, digital data is increasing due to newly developed instruments, such as conforcal laser microscopes, atomic force microscopes and so on. These save us to convert analog into digital data for a next use (e.g. image processing, analysis, recognition and reconstruction.) However, there are no data handling and processing system, which we, anatomists, are content with. Analog data, such as photographic paper, negative films and slides, also still popular in our discipline. Therefore, a combined system which handle with both analog and digital data, is nessesary. From a view point of anatomical study and “hyper-end user”, we propose an ideal integrated system for morphological study and present a prototype and a part of our three-dimensional reconstruction system in this paper.

The integrated three-dimensional reconstruction system include many steps:

1. project planning, sample preparation and observation 2. image recording, retrieving, processing, analysis and recognition 3. reconstruction, observation and re-observation of raw data 4. disscussion for the results, publishing, and exchange opinion

This could be called an integrated research environment. To realize the ideal system and environment, “the Internet” is thought to be essential. As a member of database committee of Japan Anatomical Association, I proposed a data collecting system via the Internet, which is experimentally used for collecting abstruct data (character) of the anatomical meeting. Inasmuch as this meeting is international, we would like to emphasize and introduce the Internet for promoting out study and our reseach between Japan and Korea.

We are trying to develope such an environment and improving it, now. Now, we forcus on the data recording, handling system and immage analysis. analog data and digital data,

We, anatomists, are eager to want an integrated three-dimensional reconstruction system [?] for morphological study, which means an organized integrated whole made up of diverse but interrelated and interdependent parts for researching. More concretely, the system includes not only three-dimensional reconstruction function but also other functions such as color image handling and processing, hypertext, the Internet activities [?, ?, ?] and so on. The systematized sets connected to the Internet will bring each performance into full play to facilitate our study.

Aiming at an ideal system, we are struggling to develop the system, a part of which is a color image handling system (CIH) combined with a color image processing system (CIP) [?, ?]. We applied this system to the structural analysis of mouse kidney.

Acknowledgment: We thank Mr. Tadashige Kamei and his staff (Strategic Research Div. NCI 211 Project, CANON INC.) for their support and Mr. Yoshimi Katagiri for his technical assist.